# Segmenting Action-Value Functions Over Time-Scales in SARSA via TD($\Delta$)

### v2

@@ 5 Analysis
Next, We examine our delta estimators through the lens of the bias-variance trade-off. In SARSA, bias emerges from dependence on current estimates, affecting updates over short time-scales, particularly with smaller discount factors, such as lower values of $\eta$ in $D_{m}$. Updates utilizing smaller discount factors demonstrate reduced variance due to their reliance on current estimates; however, they exhibit increased bias as they fail to integrate substantial information regarding future rewards. In contrast, variance escalates for updates associated with extended time-scales, as future reward information adds further variability. In TD($\Delta$), action-value components $D_{m}$ with higher $\eta_{m}$ are associated with greater long-term rewards; however, this results in increased variance stemming from the stochastic characteristics of reward sequences. We begin by demonstrating that our estimator is equivalent to the standard estimator $\hat{Q}_{\eta}$ under specific conditions, as outlined in Theorem 1. This comparison elucidates the essential metrics of our estimator that may indicate potential advantages over the standard $\hat{Q}_{\eta}$ estimator. Building on this result and previous research by Kearns and Singh[9], we modify these analyses for SARSA to investigate the effects of bias and variance, along with the TD($\Delta$) decomposition framework, in relation to action-value functions. Our objective is to extend the bias-variance error bound framework to action-value settings using TD($\Delta$)(i.e., Theorem 4), thereby offering a better understanding of how these quantities can be balanced to obtain optimal results[8].
@@ 5.1 SARSA Equivalency Configurations and Enhancements
In specific situations, we can show that our delta estimators for SARSA align with the conventional action-value estimator when restructured into an action-value function. This discussion centers on linear function approximation of the specified form:
@@
where $G_{n}^{\eta,\lambda}$ represents the TD($\lambda$) return defined in Eq.20. Likewise, the delta estimator approach TD($\lambda_{m},\Delta$) is used to update each $\hat{D}_{m}$:
@@
where $G_{n}^{M,\lambda_{m}}$ is specified identically to the TD($\Delta$) return defined in Eq.22, modified for the particular action-value function. In these equations, $\alpha$ and $\{\alpha_{m}\}_{m}$ represent positive learning rates. The following theorem demonstrates the equivalence of two SARSA version algorithms in this context. The subsequent theorems resemble those presented by Romoff et al. [8]; however, we offer a proof concerning the action-value function in relation to SARSA, whereas they established the proof for the value function.
@@
The proof is presented in appendix PROOFS.
@@
Inspired by Romoff et al.[8], $\forall\lambda\in\left[0,\frac{1+\eta}{2\eta}\right]$, the operator $T_{\lambda}$ is defined by the Eq. $T_{\lambda}Q=Q+(I-\lambda\eta P)^{-1}(TQ-Q)$ $,\quad\forall Q\in\mathbb{R}^{|S|\times|A|}$, is well-defined. Furthermore, $T_{\lambda}Q$ constitutes a contraction in relation to the max norm, with its contraction coefficient expressed as $\frac{\eta}{|1-\lambda\eta|}$.
The proof is presented in appendix PROOFS.
@@
Although these methods could theoretically be equivalent, such equivalence is not common. Generally, for equivalence to be maintained, the learning rate must remain consistent across all time scales. This represents a significant limitation, as shorter time scales, which involve fewer future steps, can be learned more rapidly than longer time scales that require consideration of additional steps ahead. Furthermore, in practical applications, particularly involving nonlinear function approximation such as deep neural networks, adaptive optimizers are frequently employed[30, 28]. These optimizers modify the learning rate in accordance with the complexity of the learning task, which is contingent upon the characteristics of the delta estimator and its target. The effective learning rate varies across different time scales, making it challenging to apply a uniform learning rate in this context.
Besides the learning rate, the method of dissecting the action-value function into distinct components (particularly, $D_{m}$ components) offers benefits that conventional TD learning, including non-delta estimators or elevated $D_{m}$ components, does not possess. The capacity to employ diverse k-step returns (or $\lambda$-return) across many time-scales offers improved control and flexibility in the learning process. If $k_{m}<k_{m+1},\forall m\quad(or\quad\eta_{m}\lambda_{m}<\eta_{m+1}\lambda_{m+1},\forall m)$, this approach may decrease variance while introducing bias, as shorter time-scales, which can be learned more rapidly, may not align perfectly with longer time scales in terms of the values they estimate.
@@ 5.2 Evaluation for Minimizing $k_{m}$ Values in SARSA via Phased Updates
To show that how our approach differ from the single estimator case, let’s follow the tabular phased version of k-step TD presented by kearns & Singh[9]. In SARSA, we aim to estimate the value of a state-action pair (s, a) start from each state $s\in\mathcal{S}$ and with action $a\in\mathcal{A}$, then n trajectories generated $\bigg{\{}S_{0}^{(x)}=s,a_{0},\eta_{0},...,S_{k}^{(x)},a_{k}^{(x)},\eta_{k}^{(x)},S_{k+1}^{(x)},...\bigg{\}}_{1\leq x\leq n}$ by following the policy $\pi$ and averaging over trajectories. Let $Q_{\eta,n}(s,a)$ denote the Q-value estimate at phase n for the discount factor $\eta$. For each iteration n, also called phase n, to get the phase-based estimate $Q_{\eta,n}(s,a)$ for (s, a), we average over the trajectories, resulting in:
@@
Theorem 3 below from the theorems of kearns & Singh[9] but they show it for state value function while we show an upper bound on the error for action-value function estimation defined by $\Delta_{n}^{\hat{Q}_{\eta}}=max_{s,a}\bigg{\{}\bigg{|}\hat{Q}_{\eta,n}(s,a)-Q_{\eta}(s,a)\bigg{|}\bigg{\}}$
@@
Inspired by Kearns & Singh[9], for any $10<\delta<$ 1, let $\epsilon=\sqrt{\frac{2log(2k/\delta)}{n}}$. with probability $1-\delta$,
@@
The proof is presented in appendix PROOFS.
In Eq. 28, A variance term ($\frac{1-\eta^{k}}{1-\eta}$) is arising due to sampling error from rewards collected along trajectories. In particular, $\epsilon$ bounds the deviation of the empirical average of rewards from the true expected reward. Second term ($\eta^{k}\Delta_{n-1}^{\hat{Q}_{\eta}}$) in Eq. 28 is a bias term that arises from bootstrapping off the previous phase’s estimates. Similarly, we consider a phased SARSA version of multi-step $TD(\Delta)$. For each phase n, we update each D as follows:
$\hat{D}_{m,n}(s,a)=\frac{1}{n}\sum_{x=1}^{n}\bigg{(}\sum_{i=1}^{k-1}(\eta_{m}^{i}-\eta_{m-1}^{i})r_{i}^{(x)}+(\eta_{m}^{k_{m}}-\eta_{m-1}^{k_{m}})Q_{\eta_{m-1}}(s_{n+k}^{(x)},a_{n+k}^{(x)})+\eta_{m}^{k_{m}}\hat{D}_{m}(s_{n+k}^{(x)},a_{n+k}^{(x)})$ (29)
We will now ascertain the highest limit of the phased error. The cumulative errors of each D component, denoted as $TD(\Delta)$, is expressed as $\sum_{m=0}^{M}\Delta_{n}^{m}$, where $\Delta_{n}^{m}=\max_{s,a}\bigg{\{}\bigg{|}\hat{D}_{m}(s,a)-D_{m}(s,a)\bigg{|}\bigg{\}}$. Fundamental differences between SARSA and Q-learning regarding the error bound: In SARSA, the error is ascertained by the agent’s actual actions, indicating that the Q-values are modified based on the state-action pairings encountered by the agent during exploration. The inaccuracy in Q-learning arises from the greedy action, defined as the action that maximizes the Q-value for the subsequent state, based on the premise that the agent consistently behaves optimally.
@@
Inspired by Romoff et al.[8], assume that $\eta_{0}\leq\eta_{1}\leq\eta_{2}\leq.....\leq\eta_{m}=\eta$ and $K_{0}\leq K_{1}\leq...\leq K_{m}=K$, for any $10<\delta<$ 1, let $\epsilon=\sqrt{\frac{2log(2k/\delta)}{n}}$, with probability $1-\delta$,
@@
The proof is presented in appendix PROOFS.
@@
The author examines the comparison of the limits for phased TD($\lambda$) in theorem 3 with those for phased TD($\Delta$) in theorem 4. This comparison demonstrates that phased TD($\Delta$) facilitates variance reduction equivalent to $\epsilon\sum_{m=0}^{M-1}\frac{\eta_{m}^{k_{m+1}}-\eta^{k_{m}}_{m}}{1-\eta_{m}}\leq 0$ but introduces a potential bias quantified as $\sum_{m=0}^{M-1}\bigg{(}\eta^{k_{m}}_{m}-\eta^{k_{m+1}}_{m}\bigg{)}\sum_{q=0}^{m}\Delta_{n-1}^{q}\geq 0$. Utilizing phased TD($\Delta$) to decrease $k_{m}$ synchronizes updates more effectively with recent actions, hence lowering variance from high-discounted return components. Shortened $k_{m}$ values accrue bias more swiftly, as SARSA bootstraps from current policy estimations that may vary over phases. It is observed that when all $k_{m}$ values are identical, both algorithms yield the same upper bound.
@@
In SARSA, the return estimate encompasses subsequent states and actions dictated by the policy. The expected discounted return over T steps closely approximates the infinite-horizon discounted expected return after $\mathrm{T}$, where $\mathrm{T}\approx\frac{1}{1-\eta}$[31]. Consequently, we can effectively simplify $k_{m}$ for any $\eta_{m}$ such that $k_{m}\approx\frac{1}{1-\eta_{m}}$, thereby adhering to this principle. Therefore, with $\mathrm{T}$ samples, similar to TD($\Delta$)[8], SARSA employing TD($\Delta$) utilizes $k_{m}\approx\frac{1}{1-\eta_{m}}$ to establish a suitable balance between bias and variance over all time scales significantly less than $\mathrm{T}$. By establishing $k_{m}$ for each $\eta_{m}$, where $\eta_{m}^{\frac{1}{1-\eta_{m}}}\leq\frac{1}{e}$, the approach guarantees that the effective horizon and variance are maintained within limits without requiring extensive parameter optimization. Each increase in $D_{m}$ facilitates the doubling of the effective horizon, hence ensuring logarithmic scaling with the quantity of action-value functions. SARSA can utilize this to modify additional parameters in alignment with the specified time scales, enhancing the algorithm’s efficiency in calculating long-term rewards without added complexity.
6 Experimental Setup
@@
We experimented with OpenAI Gym control tasks, and Fig.1 shows the relevant environments. The experiments were carried out on a single PC with 16 GB of memory, an Intel Core i7-2600 processor, and a GPU. The system ran 64-bit Ubuntu 22.04.5 LTS. We utilized Python 3.9 and the PyTorch package for development. We also used the OpenAI Gym toolkit, as mentioned by G. Brockman et al. [brockman2016openai]. Ten randomly selected seeds were used in each experiment, and the average outcomes are presented.
(a) FrozenLake-v1
(b) Taxi-v3
(c) Blackjack-v1
(d) CliffWalking-v0
Figure 1: OpenAI Gym control used for all experiments. In order from the left to right: Frozen Lake, Taxi, Blackjack and Cliff Walking.
@@ 6.1 Tabular
We utilize Q-value iteration to determine the optimal Q-value function. We employ an open-source version of Value Iteration (VI) derived from the code of Peter Henderson and Wei-Di Chang[value-iteration], guided by the algorithmic principles outlined in Bellman[22] and Sutton & Barto[29]. Although their methodology calculates the optimal value function, we modify it to determine the optimal Q-value function. Upon acquiring the optimal Q-values, we employ Temporal Difference (TD) and SARSA($\Delta$) to derive approximation Q-values via sampling. After 5000 steps, each learning iteration comes to an end. For every setup, we run 200 trials using different random seeds that influence the dynamics of the transition. We implement our algorithms using the code that Kostrikov provided[pytorchrl]. We compare our approach to k-step SARSA, as described in Sutton and Barto’s book, Chapter 7.2[29]. We evaluate SARSA($\Delta$) on multiple RL benchmarks, such as the Ring MDP[9] and deterministic & stochastic environments, to determine its effectiveness. A illustration of the Ring MDP is shown in Fig. 2, which shows a Markov process with five states arranged in a circle. There is a 0.95 chance of moving one state clockwise around the ring and 0.05 chance of staying in the present state at each step. While the other states on the ring offer a reward of 0, two adjacent states offer rewards of +1 and -1, respectively.
Figure 2: Five-state ring MDP as described by Kearns[9]
@@
All experiments utilize the identical parameters established by Romoff et al.[8], which involve supplying a variable number of gamma values, commencing at 0 and incrementing according to $\eta_{m+1}=\frac{\eta_{m}+1}{2}$ until the requisite maximum $\eta_{m}$ is attained. As previously mentioned in Section 5.2, $K_{m}=\frac{1}{(1-\eta_{m})}$ for every m. The baseline consists of a singular estimator with $\eta=\eta_{m}$ and $k=k_{m}$. Fig. 3(a) depicts the efficacy of two methodologies: K-step SARSA ($\Delta$) and K-step SARSA TD, employing a discount factor ($\eta$ = 0.996), k = 64 (a parameter associated with the algorithm’s step count), and diverse values of the learning rate $\alpha$. The y-axis denotes the absolute error, whereas the x-axis signifies the parameter $\alpha$, ranging from 0.0 to 1.0. At $\alpha$=0.0, both methodologies commence at an equivalent error level (about 0.00230). As $\alpha$ increases, K-step SARSA ($\Delta$) initially diminishes its inaccuracy marginally before seeing a consistent rise. K-step SARSA TD, conversely, exhibits a consistent increase in error with minimal fluctuation. Overall course: Relative to K-step SARSA TD, the K-step SARSA ($\Delta$) technique has greater stability and consistently yields superior outcomes (i.e., reduced error) across all values of $\alpha$. Both approaches exhibit sensitivity to variations in $\alpha$, although K-step SARSA TD demonstrates greater sensitivity, evidenced by its more pronounced increase in error.
@@
Fig. 3(b) demonstrates the effectiveness of the two methodologies, K-step SARSA ($\Delta$) and K-step SARSA TD, in relation to the number of k-steps. The y-axis indicates the absolute error, whereas the x-axis represents k-steps, which refers to the number of steps utilized in the algorithms. The range of k-steps extends from 0 to roughly 250. K-step SARSA ($\Delta$) commences with a reduced error at smaller k-steps, which subsequently increases with bigger k-steps. K-step SARSA TD begins with an error level comparable to K-step SARSA ($\Delta$) for low k-steps; however, its error escalates more steeply as k-steps grow. The overall pattern indicates that both techniques exhibit escalating inaccuracy as k-steps rise. K-step SARSA ($\Delta$) consistently surpasses K-step SARSA TD in all k-step values by exhibiting a reduced error rate. The swift error amplification in K-step SARSA TD indicates that the scheduling method employed in K-step SARSA ($\Delta$) mitigates error accumulation. Furthermore, in this specific MDP, the error escalates with k due to the bias-variance trade-off associated with k-step returns, aligning with the conclusions of Kearns & Singh[9, 8].
Comparison of Figs. 3(a) and 3(b): In the comparison of the two figures, K-step SARSA ($\Delta$) consistently surpasses K-step SARSA TD in both circumstances ($\alpha$ and k-steps). Both approaches exhibit sensitivity to elevated parameter values ($\alpha$ or k-steps). K-step SARSA ($\Delta$) exhibits enhanced stability and attains reduced errors as a result of its scheduling approach. Significantly, errors escalate more rapidly with K-step SARSA TD as k-steps grow in comparison to the increase in $\alpha$.
(a) Error at different learning rates
(b) Error at different k values
@@
Figure 3: Fig.3(a) illustrates the outcomes for $\eta_{M}=0.996$ on the five-state Ring MDP. The error signifies the absolute discrepancy between the estimated and actual discounted Q-value function (precomputed via Q-value iteration), averaged across the complete learning trajectory of 5000 timesteps. Fig.3(b) illustrates the average absolute error for the optimal learning rate at each k-step return, extending beyond the effective planning horizon of $\eta_{M}$.

### v4

@@ 5 Analysis
Subsequently, the delta estimators are analyzed in relation to the bias-variance trade-off. In SARSA, bias arises from reliance on current estimates, influencing updates over short time scales, especially with smaller discount factors, like lower values of $\eta$ in $D_{m}$. Utilizing smaller discount factors in updates leads to a decrease in variance because they depend on current estimates. However, this approach also results in heightened bias, as it does not incorporate significant information about future rewards. Conversely, variance increases for updates linked to longer time scales, as future reward information introduces additional variability. In TD($\Delta$), action-value components $D_{m}$ with higher $\eta_{m}$ are associated with greater long-term rewards; however, this results in increased variance stemming from the stochastic characteristics of reward sequences. We begin by demonstrating that our estimator is equivalent to the standard estimator $\hat{Q}_{\eta}$ under specific conditions, as outlined in Theorem 1. This comparison elucidates the essential metrics of our estimator that may indicate potential advantages over the standard $\hat{Q}_{\eta}$ estimator. Building on this result and previous research by Kearns and Singh [11], the analyses are adapted for SARSA to investigate the effects of bias and variance, along with the TD($\Delta$) decomposition framework, in relation to action-value functions. Our objective is to extend the bias-variance error bound framework to action-value settings using TD($\Delta$)(i.e., Theorem 4), thereby offering a better understanding of how these quantities can be balanced to obtain optimal results [1].
@@ 5.1 SARSA Equivalency Configurations and Enhancements
In certain scenarios, we can demonstrate that our delta estimators for SARSA correspond with the traditional action-value estimator when reformulated into an action-value function. This discourse focuses on the approximation of the linear function of the specified form:
@@
Here, $G_{n}^{\eta,\lambda}$ denotes the TD($\lambda$) return as stated in Eq. 19. Likewise, the delta estimator approach TD($\lambda_{m},\Delta$) is used to update each $\hat{D}_{m}$:
@@
where $G_{n}^{M,\lambda_{m}}$ is specified identically to the TD($\Delta$) return defined in Eq. 21, modified for the particular action-value function. In these equations, $\alpha$ and $\{\alpha_{m}\}_{m}$ represent positive learning rates. Two SARSA version algorithms are shown to be equivalent in this context by the following theorem. The subsequent theorems resemble those presented by Romoff et al.[1]; however, a proof is presented for the action-value function in relation to SARSA, whereas Romoff et al. established the proof for the value function.
@@
The proof is presented in appendix A.
@@
Inspired by Romoff et al.[1], $\forall\lambda\in\left[0,\frac{1+\eta}{2\eta}\right]$, the operator $T_{\lambda}$ is defined by the Eq. $T_{\lambda}Q=Q+(I-\lambda\eta P)^{-1}(TQ-Q)$ $\quad\forall Q\in\mathbb{R}^{|S|\times|A|},$ is well-defined. Furthermore, $T_{\lambda}Q$ constitutes a contraction in relation to the max norm, with its contraction coefficient expressed as $\frac{\eta}{|1-\lambda\eta|}$.
The proof is presented in appendix A.
@@
Although these methods could theoretically be equivalent, such equivalence is not common. To maintain the preservation of equivalence, it is crucial for the learning rate to remain consistent across all time scales. This underscores a significant limitation, as shorter time scales, which involve fewer future steps, can be acquired more rapidly than longer time scales that require the assessment of additional steps beforehand. Furthermore, in practical applications, particularly involving nonlinear function approximation such as deep neural networks, adaptive optimizers are frequently employed [40, 39]. The optimizers modify the learning rate accordance with the complexity of the learning task, which depends on the attributes of the delta estimator and its target. There is no universally applicable learning rate in this setting since the effective learning rate changes with time scale.
In addition to the learning rate, breaking the action-value function into separate components (especially the $D_{m}$ components) has advantages that standard TD learning, such as non-delta estimators, does not have. The ability to utilize various k-step returns (or $\lambda$-return) across multiple time scales enhances control and adaptability in the learning process. If $k_{m}<k_{m+1},\forall m\quad(or\quad\eta_{m}\lambda_{m}<\eta_{m+1}\lambda_{m+1},\forall m)$, this approach has the potential to reduce variance while simultaneously introducing bias, since shorter time scales, which can be learned more quickly, may not correspond perfectly with longer time scales regarding the values they estimate.
@@ 5.2 Evaluation for Minimizing $k_{m}$ Values in SARSA via Phased Updates
To show how our approach differs from the single estimator case, let’s follow the tabular phased version of k-step TD presented by Kearns and Singh[11]. In SARSA, the goal is to evaluate the value associated with a state-action pair (s, a) starting from each state $s\in\mathcal{S}$ and with action $a\in\mathcal{A}$, then n trajectories are generated $\bigg{\{}S_{0}^{(x)}=s,a_{0},\eta_{0},...,S_{k}^{(x)},a_{k}^{(x)},\eta_{k}^{(x)},S_{k+1}^{(x)},...\bigg{\}}_{1\leq x\leq n}$ by following the policy $\pi$ and averaging over trajectories. Let $Q_{\eta,n}(s,a)$ denote the Q-value estimate at phase n for the discount factor $\eta$. For each iteration n, also called phase n, obtain the phase-based estimate $Q_{\eta,n}(s,a)$ for (s, a); we average over the trajectories, resulting in:
@@
Theorem 3 draws upon the findings of Kearns and Singh[11], who demonstrated the result concerning the state-value function. In contrast, the subsequent theorem introduces an upper limit on the error associated with action-value function estimation, expressed as $\Delta_{n}^{\hat{Q}_{\eta}}=max_{s,a}\bigg{\{}\bigg{|}\hat{Q}_{\eta,n}(s,a)-Q_{\eta}(s,a)\bigg{|}\bigg{\}}$
@@
Inspired by Kearns and Singh[11], for any $0<\delta<1$, let $\epsilon=\sqrt{\frac{2log(2k/\delta)}{n}}$. with probability $1-\delta$,
@@
The proof is presented in appendix A.
In Eq. 27, a variance term ($\frac{1-\eta^{k}}{1-\eta}$) is arising due to sampling error from rewards collected along trajectories. In particular, $\epsilon$ bounds the deviation of the empirical average of rewards from the true expected reward. The second term ($\eta^{k}\Delta_{n-1}^{\hat{Q}_{\eta}}$) in Eq. 27 is a bias term that arises from bootstrapping off the previous phase’s estimates.
Similarly, a phased SARSA variant of the multi-step $TD(\Delta)$ method is considered. For each phase n, each D component is updated as follows:
$\hat{D}_{m,n}(s,a)=\frac{1}{n}\sum_{x=1}^{n}\biggl{(}\sum_{i=1}^{k-1}(\eta_{m}^{i}-\eta_{m-1}^{i})r_{i}^{(x)}+(\eta_{m}^{k_{m}}-\eta_{m-1}^{k_{m}})Q_{\eta_{m-1}}(s_{n+k}^{(x)},a_{n+k}^{(x)})+\eta_{m}^{k_{m}}\hat{D}_{m}(s_{n+k}^{(x)},a_{n+k}^{(x)})\biggr{)}$ (28)
The maximum threshold for the phased error has now been established. The cumulative errors of each D component, denoted as $TD(\Delta)$, is expressed as $\sum_{m=0}^{M}\Delta_{n}^{m}$, where $\Delta_{n}^{m}=\max_{s,a}\bigg{\{}\bigg{|}\hat{D}_{m}(s,a)-D_{m}(s,a)\bigg{|}\bigg{\}}$. Essential differences between SARSA and Q-learning regarding the error bound: In SARSA, the error is defined by the actions taken by the agent, indicating that the Q-values are modified based on the state-action pairings encountered by the agent during exploration. The inaccuracy in Q-learning emerges from the greedy action, which is defined as the action that maximizes the Q-value for the future state, assuming that the agent consistently operates in an optimal manner.
@@
Inspired by Romoff et al.[1], assume that $\eta_{0}\leq\eta_{1}\leq\eta_{2}\leq.....\leq\eta_{m}=\eta$ and $K_{0}\leq K_{1}\leq...\leq K_{m}=K$, for any $0<\delta<1$, let $\epsilon=\sqrt{\frac{2log(2k/\delta)}{n}}$, with probability $1-\delta$,
@@
The proof is presented in appendix A.
@@
The authors examine the comparison of the limits for phased TD($\lambda$) in theorem 3 with those for phased TD($\Delta$) in theorem 4. This comparison demonstrates that phased TD($\Delta$) facilitates variance reduction equivalent to $\epsilon\sum_{m=0}^{M-1}\frac{\eta_{m}^{k_{m+1}}-\eta^{k_{m}}_{m}}{1-\eta_{m}}\leq 0$ but introduces a potential bias quantified as $\sum_{m=0}^{M-1}\bigg{(}\eta^{k_{m}}_{m}-\eta^{k_{m+1}}_{m}\bigg{)}\sum_{q=0}^{m}\Delta_{n-1}^{q}\geq 0$. Utilizing phased TD($\Delta$) to decrease $k_{m}$ synchronizes updates more effectively with recent actions, hence lowering variance from high-discounted return components. Shortened $k_{m}$ values accrue bias more swiftly, as SARSA bootstraps from current policy estimations that may vary over phases. It is observed that when all $k_{m}$ values are identical, both algorithms yield the same upper bound.
@@
In SARSA, the return estimate includes subsequent states and actions determined by the policy. The anticipated discounted return over T steps closely approximates the infinite-horizon discounted expected return after $\mathrm{T}$, where $\mathrm{T}\approx\frac{1}{1-\eta}$ [41]. Consequently, $k_{m}$ can be effectively simplified for any $\eta_{m}$ such that $k_{m}\approx\frac{1}{1-\eta_{m}}$, thereby adhering to this principle. Therefore, with $\mathrm{T}$ samples, similar to TD($\Delta$) [1], SARSA employing TD($\Delta$) utilizes $k_{m}\approx\frac{1}{1-\eta_{m}}$ to establish a suitable balance between bias and variance over all time scales significantly less than $\mathrm{T}$. By establishing $k_{m}$ for each $\eta_{m}$, where $\eta_{m}^{\frac{1}{1-\eta_{m}}}\leq\frac{1}{e}$, the approach guarantees that the effective horizon and variance are maintained within limits without requiring extensive parameter optimization. Each increase in $D_{m}$ facilitates the doubling of the effective horizon, hence ensuring logarithmic scaling with the quantity of action-value functions. SARSA can utilize this to modify additional parameters in alignment with the specified time scales, enhancing the algorithm’s efficiency in calculating long-term rewards without added complexity.
6 Experiments and Results
Experiments were conducted in three distinct environments to evaluate the effectiveness of the proposed method: the Tabular Ring MDP environment detailed in Section 6.1, the deterministic and stochastic OpenAI Gym environments examined in Section 6.2, and the Atari environments from OpenAI Gym described in Section 6.3. This section initiates with a detailed overview of the experimental setup.
@@ 6 Experiments and Results
Experimental Setup: Experiments were conducted using OpenAI Gym control tasks [42], and Figs.1, 3, and 6 show the relevant environments. The experiments were carried out on a single PC with 16 GB of memory, an Intel Core i7-2600 processor, and a GPU. The machine ran on 64-bit Ubuntu 22.04.5 LTS. Python 3.9 was used during development, along with the PyTorch package. In each experiment, ten randomly selected seeds were utilized, and the average results are provided. The remaining hyperparameter settings are displayed in the table 4 of appendix B.
@@ 6.1 Tabular
Q-value iteration is utilized to determine the optimal Q-value function. An open-source implementation of Value Iteration (VI), based on the code by Peter Henderson[43], is employed in accordance with the algorithmic principles outlined in Bellman[33] and Sutton et al.[12]. The methodology calculates the optimal value function; however, this approach is adjusted to ascertain the optimal Q-value function. Upon acquiring the optimal Q-values, Temporal Difference (TD) and SARSA($\Delta$) methods are applied to derive approximation Q-values via sampling. After 5000 steps, each learning iteration comes to an end. For every setup, 200 trials are executed using different random seeds that influence the dynamics of the transition. The algorithms are implemented using the codebase provided by Kostrikov[44].
The proposed method is contrasted with k-step SARSA, as described in Sutton et al.[12, Chapter 7.2]. SARSA ($\Delta$) is evaluated against different RL benchmarks, including the Ring MDP [11] and both deterministic and stochastic environments, to determine its efficacy. Fig. 1 illustrates the Ring MDP, which consists of a Markov process featuring five states arranged in a circular form. The probability of transitioning one state clockwise around the ring is 0.95, while the likelihood of remaining in the current state at each step is 0.05. In contrast to the other states on the ring that provide a reward of 0, two neighboring states present rewards of +1 and -1, respectively.
Figure 1: Five-state ring MDP as described by Kearns and Singh[11]
@@
All experiments utilize the identical parameters established by Romoff et al.[1], which involve supplying a variable number of gamma values, commencing at 0 and incrementing according to $\eta_{m+1}=\frac{\eta_{m}+1}{2}$ until the requisite maximum $\eta_{m}$ is attained. As previously mentioned in Section 5.2, $K_{m}=\frac{1}{(1-\eta_{m})}$ for every m. The baseline consists of a singular estimator with $\eta=\eta_{m}$ and $k=k_{m}$. Fig. 2a illustrates the effectiveness of two methodologies: K-step SARSA ($\Delta$) and K-step SARSA TD, utilizing a discount factor ($\eta$ = 0.996), k = 64 (a parameter related to the algorithm’s step count), and various values of the learning rate $\alpha$. The y-axis represents the absolute error, while the x-axis indicates the parameter $\alpha$, which ranges from 0.0 to 1.0. At $\alpha$=0.0, both methods start with about the same amount of error (around 0.00230). As $\alpha$ increases, K-step SARSA ($\Delta$) exhibits a slight initial reduction in inaccuracy, followed by a consistent rise thereafter. In contrast, K-step SARSA TD shows a constant rise in error with low variability. Overall, the K-step SARSA ($\Delta$) approach outperforms K-step SARSA TD in terms of stability and accuracy across all $\alpha$ values. Both techniques are sensitive to fluctuations in $\alpha$, but K-step SARSA TD has a more significant rise in error.
@@
Fig. 2b compares the efficacy of two approaches, K-step SARSA ($\Delta$) and K-step SARSA TD, based on the number of k-steps. The y-axis represents the absolute error, whereas the x-axis depicts k-steps, which relates to the number of steps utilized in the algorithms. The range of k-steps spans from 0 to approximately 250. K-step SARSA ($\Delta$) begins with a lower error at smaller k-steps, which then escalates with larger k-steps. K-step SARSA TD starts with an error level similar to that of K-step SARSA ($\Delta$) for lower k-steps; however, its error increases more sharply as k-steps increase. The overall pattern indicates that both techniques exhibit a rise in inaccuracy as the number of k-steps increases. K-step SARSA ($\Delta$) consistently demonstrates superior performance compared to K-step SARSA TD across all k-step values, evidenced by a lower error rate. The rapid error amplification noted in K-step SARSA TD indicates that the scheduling method employed in K-step SARSA ($\Delta$) successfully mitigates error accumulation. Additionally, in this particular MDP, the error increases with k as a result of the bias-variance trade-off linked to k-step returns, consistent with the findings of Romoff et al.[1], Kearns and Singh[11].
Analysis of Figs. 2a and 2b: In comparing the two figures, K-step SARSA ($\Delta$) consistently outperforms K-step SARSA TD in both scenarios ($\alpha$ and k-steps). Both approaches demonstrate sensitivity to increased parameter values ($\alpha$ or k-steps). K-step SARSA ($\Delta$) demonstrates improved stability and achieves lower error rates due to its scheduling methodology. Interestingly, compared to the growth in $\alpha$, errors climb faster with K-step SARSA TD as the number of k-steps increases.
(a) Error at different learning rates.
(b) Error at different k values.
@@
Figure 2: Fig. 2a illustrates the outcomes for $\eta_{M}=0.996$ on the five-state Ring MDP. The error signifies the absolute discrepancy between the estimated and actual discounted Q-value function (precomputed via Q-value iteration), averaged across the complete learning trajectory of 5000 timesteps. Fig. 2b illustrates the average absolute error for the optimal learning rate at each k-step return, extending beyond the effective planning horizon of $\eta_{M}$.